At a point 10 m away from a long straight thin wire, the magnetic field due to the wire is 0.10 mT. What current flows through the wire? (o 4IT 10-7 T. m/A) 1.26 Submit My Answers Give Up Incorrect; Try Again; 5 attempts remaining Provide Feedback Continue
Answer #1

magnetic field due to current carrying wire

B = Uo*i/(2*pi*d)

=> 0.10*10^-3 = 4*pi*10^-7*i/(2*pi*10)

=> i = 5000 A